The crystal puller shown in figure1.3 has the following characteristics: (1) the motor speed varies as 10 rpm/volt, (2) the reducer/puller causes a rise of the vertical shaft of 0. mm/s for each 10 rpm of the motor shaft, and (3) the tach outputs 0.4 volts/rpm. a. For a steady pull rate of 1.0 cm/min, find the required motor voltage. V 1.7 and the required reference voltage, v b. Find the required ramp voltage, v,(1), if the pull up should ramp from 0.0 cm/min to 2.0 cm/min in 2.5 minutes
Answer #1

a.

Given pull rate is 0.1mm/s for each 10 rpm of motor shaft

So, for 0.1 mm/s -- 10 rpm

therefore for 1 cm/min -- \small \frac{1000}{60} rpm

and it is given that for 1 volt motor speed is 10 rpm

therefore for  \small \frac{1000}{60} rpm --   \small \frac{5}{3} volt

or Vm  = 1.67 volt (equivalent)

It is given that, tach output is 0.4 volts for 1 rpm

therefore for \small \frac{1000}{60} rpm ---- output is \small \frac{400}{60} volt or 6.67 volt(equivalent)

Error voltage Ve = Vr - Vs

Therefore Vr = Ve + Vs

The Value of Ve will be decided by design of compensator if it is required to be zero so

Vr = Vs = 6.67 volt

b.

From previous solution, required motor speed for 1 cm/min is \small \frac{1000}{60} rpm

therefore for 2 cm/min, speed would be \small \frac{2000}{60} rpm

so Tach output would be as previously, 6.67 X 2 = 13.34 volt

So for  Ve = 0

The Ramp voltage Vr(t), for change from 0.0 cm/min to 2.0 cm/min in 2.5 mins would be

\small \left ( \frac{13.34-0}{2.5} \right )t

or 5.34t (where t is in minutes)